# AGN - host galaxy photometric decomposition using a fast, accurate and precise deep learning approach

### v2

@@ 5 Summary and Conclusions
In this paper, we presented a new DL-based method Zoobot to determine the AGN contribution fraction to the total flux of a galaxy, trained on realistic mock images generated from cosmological hydrodynamical simulations. Specifically, we constructed the training sample for our DL model from the IllustrisTNG simulations. First, we mimicked the observational effects of theJWST/COSMOS-WEB survey in the NIRCAM F150W filter, by using the same pixel scale, convolving with the full range of the survey PSFs, adding Poisson noise and finally, adding the real JWST sky background. Then, we artificially added an AGN component (approximated by the PSF) at different levels to represent different AGN contribution fractions to the total flux. After training our model we analysed the performance results in detail on a test sample. Furthermore, we used GALFIT to fit the 2D light profile of the galaxies in the test sample with a combination of Sérsic + PSF profiles to calculate the AGN contribution fraction within a circular aperture. We showed how our DL method significantly outperforms the traditional 2D surface brightness fitting method, in precision, accuracy, speed and success rate. Our main results are summarised in the following:
@@
    The AGN contribution fraction $f_{AGN}$ predicted by Zoobot has a very tight correlation with the injected $f_{AGN}$, with a mean difference of $-0.0018$ and an overall RMSE $=0.013$ for the whole test sample over the redshift range $30.5<z<$ 3, demonstrating how well our DL method is at recovering the true AGN contribution to the total flux. The RMSE performance metric is around 10 times better than that obtained using GALFIT. In both methods, there is an increase in the RMSE at lower $f_{AGN}$ values, higher redshifts, lower S/N and smaller galaxy sizes.
@@
    The outlier fraction given by Zoobot, as measured by the number of galaxies that have RAE $>20\%$, is below 8% for most of the sample, except for galaxies with very low AGN contribution fractions ($f_{AGN}<0.1$) and faint galaxies (mag $>22$ in the F150W filter). There is only a small increase with lower S/N. However, the outlier fraction is most affected by the AGN contribution fractions and starts to increase sharply at $f_{AGN}<0.2$. The outlier fraction obtained from the GALFIT results is on average around 20%, and similarly to our Zoobot results, there is an increase with decreasing $f_{AGN}$ and S/N. Furthermore, there is a significant increase in the outlier fraction at higher redshift, which is not observed in the Zoobot results.

### v3

@@ 5 Summary and Conclusions
In this paper, we presented a new DL-based method Zoobot to determine the AGN contribution fraction to the total flux of a galaxy, trained on realistic mock images generated from cosmological hydrodynamical simulations. Specifically, we constructed the training sample for our DL model from the IllustrisTNG simulations. First, we mimicked the observational effects of theJWST/COSMOS-WEB survey in the NIRCAM F150W filter, by using the same pixel scale, convolving with the full range of the survey PSFs, adding Poisson noise and finally, adding the real JWST sky background. Then, we artificially added an AGN component (approximated by the PSF) at different levels to represent different AGN contribution fractions to the total flux. After training our model we analysed the performance results in detail on a test sample. Furthermore, we used GALFIT to fit the 2D light profile of the galaxies in the test sample with a combination of Sérsic + PSF profiles to calculate the AGN contribution fraction within a circular aperture. We showed how our DL method outperforms the traditional 2D surface brightness fitting method and is 2500 times faster in inferring the AGN fraction. Our main results are summarised in the following:
@@
    The AGN contribution fraction $f_{AGN}$ predicted by Zoobot has a very tight correlation with the injected $f_{AGN}$, with a mean difference of 0.0018-0.0018 and an overall RMSE =0.013=0.013 for the whole test sample over the redshift range $0.5<z<3$, demonstrating how well our DL method is at recovering the true AGN contribution to the total flux. The RMSE performance metric is around 10 times better than that obtained using GALFIT. In both methods, there is an increase in the RMSE at lower $f_{AGN}$ values, higher redshifts, lower S/N and smaller galaxy sizes.
@@
    The outlier fraction given by Zoobot, as measured by the number of galaxies that have RAE $20\%$, is below 8% for most of the sample, except for galaxies with very low AGN contribution fractions ($f_{AGN}<0.1$) and faint galaxies (mag >22>22 in the F150W filter). There is only a small increase with lower S/N. However, the outlier fraction is most affected by the AGN contribution fractions and starts to increase sharply at $f_{AGN}<0.2$. The outlier fraction obtained from the GALFIT results is on average around 20%, and similarly to our Zoobot results, there is an increase with decreasing $f_{AGN}$ and S/N. Furthermore, there is a significant increase in the outlier fraction at higher redshift, which is not observed in the Zoobot results.
@@ Appendix B AGN example images
Appendix C Color-color selection
C.1 Training in Multiple JWST Filters
We retrained our deep learning model on mock data constructed for the additional NIRCam filters: F115W, F277W, and F [^4] W. The mock data for these filters were generated analogously to F150W (see Sect. 2.3), with the only differences being the use of the specific filter response curves and PSF models for each band, and the use of cutouts corresponding to the respective filter. The procedure for creating model AGN-host galaxies is identical to that described for F150W (see Sect. 3.1). The training procedure followed the same setup as for F150W (see Sect. 3.2 of the main text), employing the same ConvNeXt-Base architecture pretrained on Zoobot for each dataset.
Table 5 summarises the model performance in each filter, reporting the RMSE, RAE, mean difference, and outlier fractions (at 20% and 30% thresholds) in the recovery of $f_{AGN}$ across the test sets. The performance is comparable across all filters, indicating that the model is robust to band choice.
Table 2: Zoobot model performance across JWST filters. [^5] We summarise the overall performance from Zoobot, in terms of the RMSE, RAE, mean difference and outlier fractions (at different percentage levels, 20% and 30%).
RMSE RAE $\langle{\Delta f_{AGN}}\rangle$ Outlier (20%) Outlier (30%)
F150W $013\pm 0.005 0.0$$76\pm 0.005 0.0$$018\pm 0.0002 0.0$$65\pm 0.004 0.0$$061\pm 0.003$
F115W $0.046\pm 0.006$ $1.14\pm 0.04$ $0.0033\pm 0.0002$ $147\pm 0.004 0.0$$084\pm 0.003$
F277W $0.026\pm 0.002$ $16\pm 0.02 0.0$$0034\pm 0.0001$ $110\pm 0.010 0.0$$079\pm 0.010$
F [^4] W $0.029\pm 0.002$ $83\pm 0.03 0.0$$0047\pm 0.0001$ $0.140\pm 0.004$ $0.104\pm 0.004$
C.2 Application to Real JWST Data and AGN Flux Calculation
We applied the trained models to the real JWST/NIRCam data in each filter (F115W, F150W, F277W, and F [^4] W) for the same galaxy sample described in the main text, obtaining a prediction of $f_{AGN}$ in each band. To compute the AGN fluxes, we require total galaxy fluxes, which we adopt from Chen et al. (2025, in prep). Briefly, Chen et al. (2025, in prep) used JWST/NIRCam images reduced by Zhuang et al. (2024). The images were processed using JWST pipeline v1.10.2 with CRDS v11.17.0, along with additional custom steps. Source photometry and morphology were derived by modelling each galaxy with a single Sérsic profile using GALFIT (Peng et al., 2002), with initial parameters from SExtractor (Bertin & Arnouts, 1996). This process provides both geometric (e.g., axis ratio, position angle) and physical (e.g., flux, magnitude, Sérsic index, effective radius) properties.
The resulting JWST/NIRCam magnitudes were cross-validated against the nearest bands in COSMOS2020 (F115W vs. VISTA NB118, F150W vs. VISTA H, F277W vs. VISTA Ks, and F [^4] W vs. IRAC CH2), showing tight correlations and good agreement. The total fluxes are then combined with the $f_{AGN}$ values from our models to derive the nuclear (AGN) fluxes in each band.
C.3 Multi-Band Analysis
Using the AGN fractions derived in each JWST/NIRCam filter, we performed a multi-band analysis to assess the nature of the central light excess, by constructing color–color diagrams in various filter combinations (e.g., F150W–F277W vs. F115W–F150W). In Figure 20 we show a subsample of colour–colour diagrams constructed from the total galaxy fluxes in the four JWST/NIRCam filters.
[…]
Figure 20: Colour–colour diagrams for eight representative filter combinations from the four JWST/NIRCam bands (F115W, F150W, F277W, F [^4] W). Black contours show the distribution of non-AGN galaxies ($f_{AGN}<0.05$ in all bands). Orange contours indicate the $f_{AGN}$-selected AGN based on the whole-galaxy fluxes. MIR-selected AGN are overplotted as red crosses (if they satisfy $f_{AGN}>5\sigma$).
Figure 21: Normalised colour distributions of our galaxy sample (F115W-F227W, F115W-F [^4] W, and F150W-F [^4] W). Galaxies with negligible nuclear light ($f_{AGN}<0.05$ in all filters and not selected as MIR AGN) are shown in black, AGN candidates with significant nuclear light ($f_{AGN}>5\sigma$ in all bands) in orange, and MIR-selected AGN in red. Vertical lines indicate the median magnitude of each distribution.
By performing the $f_{AGN}$ analysis across multiple JWST/NIRCam filters, we can partially mitigate the rest-frame wavelength differences introduced by the wide redshift range of our sample ($0.5<z<3$). While any single filter probes different physical regimes at different redshifts, the combination of multiple bands allows us to compare nuclear light excesses more consistently across the sample. This approach reduces, though does not fully eliminate, potential biases due to varying SED coverage. A more detailed analysis accounting explicitly for rest-frame band differences and nuclear versus host SEDs will be required to fully quantify these effects.